1. Suppose that the firm operates in a perfectly competitive market. The market price of its product is 4 Birr. The firm estimates its cost of production with the following cost function:

TC=500+20q-5q2+0.33q3

a. What level of output should the firm produce to maximize its profit?

b. Determine the level of profit at equilibrium.

c. What minimum price is required by the firm to stay in the market?

2. Consider the following information for a particular economy.

  • Total population=60 million
  • Total labour force=40 million
  • Number of employed=30 million.

Natural rate of unemployment=12%

a. Find the total unemployment rate.

b. Calculate the cyclical unemployment.


1
Expert's answer
2022-01-02T18:21:54-0500
  1. a. P = MR = MC

"MC= (5+20q-5q^2+0.33q^3)'=20-10q+0.99q^2"

"20-10q+0.99q^2=4"

"0.99q^2-10q+16=0"

q=2

q=8


b."Profit=TR - TC=4\times8-(5+20\times8-5\times8^2+0.33\times8^3)=18,04"

c. P = AVC = VC/Q

"\frac{5+20q-5q^2+0.33q^3}{q}=\frac{5}{q}+20-5q+0.33q^2"

q=8

"P=\frac{5}{q}+20-5q+0.33q^2=\frac{5}{8}+20-5\times8+0.33\times8^2=1.745"

2.

"u=\frac{U}{L}=\frac{30\times0.12}{30}=0.107=0.12"

uc=u-u*=0.12-0.12=0